OpenStudy (anonymous):

PLEASE HELP! Find b and c so that y=5x^2+bx+c has vertex (9,8).

OpenStudy (anonymous):

hint: (h,k) = where h = -b/(2a) k = f(h)

OpenStudy (anonymous):

go to the site quickmath.com

hero (hero):

\(x = -\dfrac{b}{2a}\) Plug in the given values of x = 9 a = 5 Then solve for b

hero (hero):

Then plug in x = 9, y = 8, and the value of b you found into the original formula and solve for c.

hero (hero):

You're welcome

OpenStudy (anonymous):

You are my hero @Hero

hero (hero):

:)